Grande03:

Hamid published a book of poems. He sold 35 copies in the first week for $10 per copy. He wanted to increase the price of each book to increase his revenue. He found out that for every 50-cent increase, x, in the price per copy, the average number of books he sold decreased by 2 per week. Use this information to create expressions that model the number of books he sold and the revenue he got. The number of books sold is given by the expression ______, and the revenue is given by the expression ______.

YasssPoo:

y = -x^2 − 2.5x + 350 Step-by-step explanation: y = (10 + 0.5x)(35 − 2x) y = 350 − 20x + 17.5x − x^2 y = 350 − 2.5x − x^2

YasssPoo:

medal plz nd tnx

Grande03:

theres two blanks and you gave me one equation